# Communication-Efficient Wireless Federated Fine-Tuning for Large-Scale AI Models

### v1

@@ Abstract
Transformer-based large language models (LLMs) have achieved remarkable success across various tasks. Yet, fine-tuning such massive models in federated learning (FL) settings poses significant challenges due to resource constraints and communication overhead. Low-Rank Adaptation (LoRA) addresses these issues by training compact, low-rank matrices instead of fully fine-tuning large models. This paper introduces a wireless federated LoRA fine-tuning framework that optimizes both learning performance and communication efficiency. We provide a novel convergence analysis, revealing how LoRA rank and covariance effects influence FL training dynamics. Leveraging these insights, we propose Sparsified Orthogonal Fine-Tuning (SOFT), an adaptive sparsification method that streamlines parameter updates without expensive matrix multiplications and singular value decomposition (SVD) operations. Additionally, we present a Two Stage Federated Algorithm (TSFA) that pre-determines LoRA rank offline and dynamically adjusts other parameters online, ensuring efficient training under latency constraints. Our framework thus enables scalable, resource-efficient deployment of large models in wireless FL scenarios.
@@ I Introduction
As a promising solution , parameter-efficient fine-tuning (PEFT) methods have been developed [2, 3]. PEFT techniques focus on adjusting a smaller subset of model parameters, thereby reducing computational load and memory requirements while achieving performance comparable to that of fully fine-tuned models. Among these methods, Low-Rank Adaptation (LoRA) [3] stands out as a particularly effective approach. LoRA minimizes the number of trainable parameters by introducing trainable low-rank matrices integrated with the existing pre-trained model weights, which remain frozen during fine-tuning. This technique allows for efficient fine-tuning by updating only these additional low-rank matrices, significantly reducing computational and memory demands without substantially compromising performance. On many tasks, LoRA has been shown to achieve comparable or even superior performance compared to fully fine-tuning. As a result, LoRA facilitates the deployment of advanced LLMs in environments with limited resources, enhancing the practicality and accessibility of sophisticated capabilities.
@@
In many practical scenarios, the data necessary for fine-tuning is distributed across multiple clients, and centralizing this data raises significant privacy and security concerns. To address these challenges, Federated Learning (FL) has emerged as a promising paradigm for collaborative model training without the need for centralized data aggregation [4, 5, 6, 7]. FL enables multiple clients to jointly train a global model by sharing model updates. Each client trains the model locally using its own data and periodically communicates the model parameters or gradients to a central server, which aggregates them to update the global model. In practical wireless FL systems, however, communication efficiency becomes critical due to limited bandwidth and fluctuating channel conditions [8]. To mitigate these challenges, sparsification techniques are commonly employed in FL by transmitting only the most significant model updates, typically by selecting the top-$k$ norm elements [9].
@@
Several recent studies have attempted to combine LoRA with FL to harness the benefits of both methodologies [10, 11, 12]. For instance, the authors in [10] pioneered the integration of LoRA-based local updates with the Federated Averaging (FedAvg) algorithm [4] for model aggregation. Building upon this, [11] proposed a method for handling heterogeneous LoRA ranks among clients, suggesting that the server zero-pads the missing components corresponding to absent rank indices in clients with smaller ranks so that all matrices match the largest rank prior to aggregation. Meanwhile, [12] introduced Federated Freeze A LoRA (FFA-LoRA), a method that addresses data heterogeneity and provides privacy guarantees through differential privacy. FFA-LoRA uses only one of the two LoRA matrices during training and employs the Gaussian mechanism to analyze differential privacy. Finally, the authors in [13] investigated the impact of data heterogeneity on learning efficiency and introduced a data-driven initialization method to address this challenge.
Despite these advancements, they share a certain limitation. All the aforementioned works rely on a determined rank and do not provide a systematic method for selecting an optimal rank that balances performance and resource constraints. Since the performance of LoRA is highly sensitive to the choice of rank $r$ [3], this oversight leads to suboptimal model performance or unnecessary resource consumption. A rank that is too low may fail to capture essential features, leading to suboptimal model performance, while a higher rank makes the clients communicate larger model sizes with the server.
@@
Moreover, a separate challenge arises when integrating sparsification methods in LoRA-based FL. In LoRA, model updates are factored into two low-rank matrices. Naively applying top-$k$ selection to each matrix independently disregards the fact that small entries in one matrix can have a disproportionately larger influence than other entries when multiplied by large entries in the other. To accurately capture the most influential updates, one would need to compute the full product of the two matrices, then apply a singular value decomposition (SVD) and transmit only the components corresponding to the largest singular values. However, frequent SVD operations are computationally prohibitive, especially for resource-constrained clients, making both naive top-$k$ and SVD-based approaches infeasible in LoRA-based FL. This underscores the need for novel sparsification methods that can select significant information while keeping computational costs feasible.
@@
To address these challenges, we propose a practical wireless federated LoRA fine-tuning framework tailored for resource-constrained clients. Our approach dynamically optimizes the LoRA rank to balance learning performance with communication efficiency, and introduces a novel LoRA sparsification strategy, specifically not requiring full matrix multiplications and SVD operations. We specifically propose Sparsified Orthogonal Fine-Tuning (SOFT), which leverages the orthogonality constraints on LoRA matrices. In this scheme, each rank vector—comprising the corresponding column vectors in the left LoRA matrix and row vectors in the right LoRA matrix—is orthogonal. As a result, the first LoRA matrix effectively functions similarly to the left singular matrix in an SVD, while the second corresponds to the right singular matrix. To determine the importance of the $i$-th rank vector, we compute the product of the norm of its column in the left matrix and the norm of its row in the right matrix, using this result as a proxy for the singular value. By enforcing these orthogonality constraints, SOFT efficiently select the most significant updates while maintaining low computational overhead.
@@
On the other hand, to determine LoRA rank, we perform a convergence analysis of LoRA within the FL setting. By incorporating LoRA-rank effects into our analysis, we derive theoretical insights into how this parameter affects the model performance and convergence speed. Based on this analysis, we first divide the optimization parameters into two groups: parameters that must be determined before the training begins and parameters that can be adjusted dynamically during the training. Specifically, since the LoRA rank represents the model structure used in FL training which must be established prior to the training, we determine it through an offline stage. After training commences, we adaptively adjust the sparsification ratios and bandwidth allocation strategy in response to real-time network conditions and resource availability in an online stage. To manage long-term latency constraints and ensure stable convergence, we employ Lyapunov optimization techniques.
@@
  • We introduce a new LoRA sparsification technique, SOFT, that avoids expensive computational operations by enforcing orthogonality in the LoRA matrices. Through an orthogonal regularization term in the loss function, SOFT ensures that these rank matrices behave similarly to left and right singular matrices in an SVD.
  • We conduct theoretical convergence analysis that explicitly incorporates the LoRA rank into FL for the first time. This analysis offers insights into how rank selection impacts model accuracy, sparsification error, and convergence speed.
  • We propose a Two Stage Federated Algorithm (TSFA) framework that first determines the LoRA rank under approximate channel conditions before the training begins, and then adaptively adjusts the sparsification ratio and bandwidth allocation at each iteration using Lyapunov optimization.
  • We rigorously validate the effectiveness of the proposed framework through extensive simulations and experiments on benchmark datasets. The results demonstrate that our approach achieves comparable or superior performance to the existing methods while significantly reducing communication overhead, thereby substantiating its practicality and efficiency in real-world FL applications.
@@ II-A Low-rank Adaptation
Rather than updating $\Delta\theta$ directly, LoRA seeks to represent the weight updates as a low-rank decomposition $\Delta\theta=\theta_{B}\theta_{A}$, where $\theta_{B}\in\mathbb{R}^{d\times r}$, $\theta_{A}\in\mathbb{R}^{r\times\ell}$, and $r\ll\min(d,\ell)$. In this way, the output can be represented as $\boldsymbol{y}=\boldsymbol{x}\theta_{P}+\frac{\alpha}{r}\boldsymbol{x}\Delta\theta=\boldsymbol{x}\theta_{P}+\frac{\alpha}{r}\boldsymbol{x}\theta_{B}\theta_{A},$ where $\boldsymbol{y}\in\mathbb{R}^{\ell\times 1}$, $\boldsymbol{x}\in\mathbb{R}^{d\times 1}$, and $\frac{\alpha}{r}$ is a scaling parameter. In LoRA, the pre-trained weights $\theta_{P}$ are kept fixed during fine-tuning, and only the low-rank matrices $\theta_{A}$ and $\theta_{B}$ are updated. This significantly reduces the number of trainable parameters by a factor of $\mathcal{O}\left(\frac{r}{\min(d,\ell)}\right)$ compared to full fine-tuning.
@@ II-B Federated LoRA Fine-Tuning
Each client $k\in\mathcal{N}$ owns a local dataset $\mathcal{D}_{k}=\{(x_{i},y_{i})\}_{i=1}^{D_{k}}$ with $D_{k}=|\mathcal{D}_{k}|$ data pairs such that $\mathcal{D}=\cup_{k\in\mathcal{N}}\mathcal{D}_{k}$. The local loss function of client $k$ is $F_{k}(\theta_{B},\theta_{A})=\frac{1}{D_{k}}\sum_{x_{i}\in\mathcal{D}_{k}}f(\theta_{B},\theta_{A};x_{i},y_{i}),$ where $f(\theta_{B},\theta_{A})$ represents the sample-wise loss function. The global loss function is $F(\theta_{B},\theta_{A})=\sum_{k\in\mathcal{K}^{t}}p_{k}F_{k}(\theta_{B},\theta_{A}),$ where $p_{k}=|\mathcal{D}_{k}|/|\mathcal{D}|$. To achieve a global model $\theta^{*}=\theta_{P}+\frac{\alpha}{r}\theta_{B}^{*}\theta_{A}^{*}$, clients periodically send updates to the server, assuming no direct interaction among the clients. The optimal global model $\theta^{*}$ for LoRA is identified through the following minimization process $\theta^{*}\!=\!\underset{\theta_{B}\in\mathbb{R}^{d\times r},\theta_{A}\in\mathbb{R}^{r\times\ell}}{\arg\min}F(\theta_{B},\theta_{A}).$ Within the wireless federated LoRA fine-tuning framework, the objective function $F(\theta_{B},\theta_{A})$ is progressively minimized through a series of local and global updates.
@@
    Global Matrices Broadcasting: The server distributes the $t$-th round global matrices $\theta_{B}^{t}$ and $\theta_{A}^{t}$ to the set of scheduled clients $\mathcal{K}^{t}$ by broadcasting. Since the server has enough power to transmit the global matrices correctly, we assume the downlink transmission is error-free.
@@
    Local Fine-Tuning: Each client $k\in\mathcal{K}^{t}$ initialize the global matrices as an initial point of local matrices, i.e. $\theta_{B,k}^{t}=\theta_{B}^{t},\theta_{A,k}^{t}=\theta_{A}^{t}$. Each client performs local training through gradient descent or ADAM resulting in an updated local models $\theta_{B,k}^{t+1}$ and $\theta_{A,k}^{t+1}$.
@@
where $p_{k}=|\mathcal{D}_{k}|/|\mathcal{D}|$ denotes the proportion of the dataset associated with client $k$, i.e. $\sum_{k\in\mathcal{N}}p_{k}=1$. After a total of $T$ rounds of this process, the training completes, yielding an optimally trained global model for LoRA. Although a client may intend to send the complete matrix product $\theta_{B,k}\theta_{A,k}$, to avoid the computational burden of matrix multiplication at the resource-constrained clients, it instead sends $\theta_{B,k}$ and $\theta_{A,k}$. This is because transmitting the full product would involve sending a matrix of size $d\times\ell$, which is generally much larger than $\theta_{B,k}$ and $\theta_{A,k}$ due to $r\ll\min\{d,\ell\}$. Therefore sending the two low-rank matrices not only minimizes local computation but also reduces the overall communication latency and overhead, as the total number of transmitted parameters is significantly low. Fig. 1 provides an overall framework.
@@ II-C Communication Model
We consider a wireless multiple access system, specifically a frequency division multiple access (FDMA) scheme for local model transmission, with a total available bandwidth of $B$. The channel gain between the server and client is modeled as $h_{k}^{t}=s_{k}^{t}q_{k}^{-\gamma}$, where $s_{k}^{t}\sim\mathcal{N}(0,1)$ represents the small-scale channel fading coefficient, and $q_{k}^{-\gamma}$ denotes the distance-dependent path loss with exponent $\gamma$. Additionally, $b_{k}^{t}\in[0,1]$ represents the bandwidth allocation ratio for client $k\in\mathcal{N}$ where $\sum_{k=1}^{N}b_{k}^{t}=1$. The variable $a_{k}^{t}\in\{0,1\}$ indicates whether client $k$ is selected in iteration $t$, where $\sum_{k=1}^{N}a_{k}^{t}=K^{t}$ and $a_{k}^{t}=0$ implies no bandwidth is allocated to client $k$, resulting in $b_{k}^{t}=0$. The transmission rate for client $k$ can be expressed as $r_{k}^{t}=b_{k}^{t}B\log_{2}\left(1+\frac{|h_{k}^{t}|^{2}}{\sigma^{2}}\right)$, where $\sigma^{2}$ represents the variance of the additive white Gaussian noise (AWGN). During the uplink phase, we assume that the amount of information per model parameter is denoted by $v$. Thus, the communication delay for client $k$ at iteration $t$ can be written as $D_{k}^{t}=\frac{a_{k}^{t}vr(d+\ell)}{r_{k}^{t}}=\frac{a_{k}^{t}vr(d+\ell)}{b_{k}^{t}B\log_{2}\left(1+\frac{|h_{k}^{t}|^{2}}{\sigma^{2}}\right)}.$ Note that if $a_{k}^{t}=0$, then $D_{k}^{t}=0$. Since the overall transmission delay at iteration $t$ is determined by the slowest client, it is given by $D^{t}=\max_{k\in\mathcal{K}^{t}}D_{k}^{t}.$

### v2

# Two-Stage Wireless Federated LoRA Fine-Tuning with Sparsified Orthogonal Updates

@@ Abstract
Federated fine-tuning of pre-trained models with Low-Rank Adaptation (LoRA) has emerged as a promising approach for privacy-preserving on-device adaptation. However, in wireless federated settings, existing methods still face two major challenges: selecting an appropriate structural LoRA rank under communication constraints and sparsifying LoRA updates without destroying the inherent structure between the two LoRA factors. In particular, naively applying conventional sparsification to the two LoRA factors incurs severe performance loss, while repeated full matrix multiplication and singular value decomposition are computationally prohibitive for resource-constrained clients. To address these issues, we propose a wireless federated LoRA fine-tuning framework with Sparsified Orthogonal Fine-Tuning (SOFT) and a Two-Stage Federated Algorithm (TSFA). SOFT imposes near orthogonality on the LoRA factors and uses a per-rank norm-based importance metric to perform efficient SVD-free sparsification. TSFA first determines the structural LoRA rank in an offline stage and then adaptively optimizes the sparsification ratio and bandwidth allocation online via Lyapunov optimization under long-term latency constraints. We further present a convergence analysis that explicitly characterizes the effects of structural LoRA rank, sparsification error, and the covariance induced by separate aggregation of LoRA factors.
@@ I Introduction
As a promising solution to fine-tuning, parameter-efficient fine-tuning (PEFT) methods have been developed [2, 3]. PEFT techniques focus on adjusting a smaller subset of model parameters, thereby reducing computational load and memory requirements while achieving performance comparable to that of fully fine-tuned models. Among these methods, Low-Rank Adaptation (LoRA) [3] stands out as a particularly effective approach. LoRA minimizes the number of trainable parameters by introducing trainable low-rank matrices integrated with the existing pre-trained model weights, which remain frozen during fine-tuning. This technique allows for efficient fine-tuning by updating only these additional low-rank matrices, significantly reducing computational and memory demands without substantially compromising performance. On many tasks, LoRA has been shown to achieve comparable or even superior performance compared to fully fine-tuning.
@@
In many practical scenarios, the data necessary for fine-tuning is distributed across multiple clients, and centralizing this data raises significant privacy and security concerns. To address these challenges, federated learning (FL) has emerged as a promising paradigm for collaborative model training without the need for centralized data aggregation [4, 5]. FL enables multiple clients to jointly train a global model by sharing model updates. Each client trains the model locally using its own data and periodically communicates the model parameters or gradients to a central server, which aggregates them to update the global model. In practical wireless FL systems, however, communication efficiency becomes critical due to limited bandwidth and fluctuating channel conditions [6, 7]. To mitigate these challenges, sparsification techniques are commonly employed in FL by transmitting only the most significant model updates, typically by selecting the top-$q$ norm elements [8].
@@
Several recent studies have attempted to combine LoRA with FL to harness the benefits of both methodologies [9, 10, 11, 12, 13, 14, 15]. For instance, the authors in [9] pioneered the integration of LoRA-based local updates with the Federated Averaging (FedAvg) algorithm [4] for model aggregation. Building upon this, [10] introduced Federated Freeze A LoRA (FFA-LoRA), a method that addresses data heterogeneity and provides privacy guarantees through differential privacy. FFA-LoRA uses only one of the two LoRA matrices during training and employs the Gaussian mechanism to analyze differential privacy. The authors in [11] investigated the impact of data heterogeneity on learning efficiency and introduced a data-driven initialization method to address this challenge. In addition, [12] proposed a personalized LoRA aggregation framework in which each client adaptively fuses the global and local LoRA parameters element-wise before local training, improving performance under heterogeneous data distributions. The authors in [13, 14] proposed a split federated LoRA framework, where the computationally intensive encoder is deployed at the edge server, while the embedding and task modules remain at the clients to account for their limited computational capabilities. Meanwhile, [15] combined LoRA with dropout-based aggregation to mitigate overfitting during training.
Despite these advances, an important problem remains unresolved. Most existing studies fix a pre-defined LoRA rank prior to training and do not offer a principled mechanism to select it under performance and resource constraints. Although some prior works [16, 17, 18] consider adaptive rank strategies, they typically operate within a pre-defined rank budget rather than determining the structural LoRA rank itself in a principled and communication-aware manner. Since performance is highly sensitive to the choice of rank r\!r\![3], this limitation can lead to either degraded model performance or inefficient resource utilization. Specifically, a rank that is too low may fail to capture essential features, whereas a higher rank increases the communication payload between clients and the server.
@@
Moreover, most existing methods do not explicitly address time varying wireless communication conditions. Although some prior works [19, 18] attepted to reduce communication overhead through top-$q$ sparsification and uniform quantization, they applied these techniques directly to LoRA without considering its structural properties, where each update is represented by two coupled low-rank matrices. Specifically, in LoRA, model updates are factorized into two low-rank matrices. Naively applying top-$q$ selection to each matrix independently ignores the interdependence between them: small entries in one matrix may exert a disproportionately large influence when multiplied by large entries in the other. Accurately identifying the most influential updates would therefore require computing the full matrix product, followed by singular value decomposition (SVD), and transmitting only the components associated with the largest singular values. However, performing frequent SVD operations is computationally prohibitive, particularly for resource-constrained clients. As a result, both naive top-$q$ approaches and SVD-based methods become impractical in LoRA-based FL. This limitation highlights the need for novel sparsification strategies that can effectively capture the most significant information while maintaining computational efficiency.
@@
To address these challenges, we propose a practical wireless federated LoRA fine-tuning framework tailored for resource-constrained clients. Our framework first selects a structural LoRA rank offline in a principled way and then adapts the sparsification ratio and bandwidth allocation online during training. Also, we introduce a novel LoRA sparsification strategy, sparsified orthogonal fine-tuning (SOFT), which leverages near orthogonality constraints on the LoRA matrices, without requiring full matrix multiplications or SVD operations. In this scheme, each rank vector, comprising the corresponding column vectors in the left LoRA matrix and row vectors in the right LoRA matrix, is near orthogonal. As a result, the first LoRA matrix effectively functions similarly to the left singular matrix in an SVD, while the second corresponds to the right singular matrix. To determine the importance of the $i\in\{1,...,r\}$-th rank vector, we compute the product of the norm of its column in the left matrix and the norm of its row in the right matrix, using this result as a proxy for the singular value. By enforcing these near orthogonality constraints, SOFT efficiently selects the most significant updates while maintaining low computational overhead.
@@
On the other hand, to determine LoRA rank, we perform a convergence analysis of LoRA within the FL setting. By incorporating the effects of structural LoRA rank, sparsification, and the discrepancy induced by separate aggregation of the two LoRA factors, we derive theoretical insights into their impact on model performance and convergence behavior. Based on this analysis, we first divide the optimization parameters into two groups: parameters that must be determined before the training begins and parameters that can be adjusted dynamically during the training. Specifically, since the LoRA rank represents the model structure used in FL training which must be established prior to the training, we determine it through an offline stage. After training commences, we adaptively adjust the sparsification ratios and bandwidth allocation strategy in response to real-time network conditions and resource availability in an online stage. To manage long-term latency constraints and ensure stable convergence, we employ Lyapunov optimization.
@@
  • We develop SOFT, a LoRA sparsification method that encourages near orthogonality in the LoRA matrices via orthogonal regularization and employs a per-rank norm-based score as an SVD-free importance proxy. This enables efficient update selection without explicitly forming full product matrix or repeatedly performing SVD.
  • We provide a convergence analysis for federated LoRA fine-tuning that explicitly captures the effects of the structural LoRA rank, sparsification error, and the discrepancy induced by separate aggregation of the two LoRA factors. The analysis clarifies how structural rank selection influences approximation quality, communication efficiency, and convergence behavior.
  • We propose a two-stage federated algorithm (TSFA) framework that first determines the structural LoRA rank under approximate channel conditions before the training begins since it decides the model structure, and then adaptively adjusts the sparsification ratio and bandwidth allocation at each iteration using Lyapunov optimization.
  • We rigorously validate the effectiveness of the proposed framework through extensive simulations and experiments on benchmark datasets. The results demonstrate that our approach achieves comparable or superior performance to the existing methods while significantly reducing communication overhead, thereby substantiating its practicality and efficiency.
@@ II-A Low-rank Adaptation
Rather than updating $\Delta\theta$ directly, LoRA seeks to represent the weight updates as a low-rank decomposition $\Delta\theta=\theta_{B}\theta_{A}$, where $\theta_{B}\in\mathbb{R}^{d\times r}$, $\theta_{A}\in\mathbb{R}^{r\times\ell}$, and $r\ll\min(d,\ell)$. In this way, the output can be represented as
$\boldsymbol{y}=\boldsymbol{x}\theta_{P}+\frac{\alpha}{r}\boldsymbol{x}\Delta\theta=\boldsymbol{x}\theta_{P}+\frac{\alpha}{r}\boldsymbol{x}\theta_{B}\theta_{A},$ (1)
where $\boldsymbol{y}\in\mathbb{R}^{\ell\times 1}$, $\boldsymbol{x}\in\mathbb{R}^{d\times 1}$, and $\frac{\alpha}{r}$ is a scaling parameter. In LoRA, the pre-trained weights $\theta_{P}$ are kept fixed during fine-tuning, and only the low-rank matrices $\theta_{A}$ and $\theta_{B}$ are updated. This significantly reduces the number of trainable parameters by a factor of $\mathcal{O}\left(\frac{r}{\min(d,\ell)}\right)$ compared to full fine-tuning.
@@ II-B Federated LoRA Fine-Tuning
Each client $k\in\mathcal{N}$ owns a local dataset $\mathcal{D}_{k}=\{(x_{i},y_{i})\}_{i=1}^{D_{k}}$ with $D_{k}=|\mathcal{D}_{k}|$ data pairs such that $\mathcal{D}=\cup_{k\in\mathcal{N}}\mathcal{D}_{k}$. The local loss function of client $k$ is
$F_{k}(\theta_{B},\theta_{A})=\frac{1}{D_{k}}\sum_{x_{i}\in\mathcal{D}_{k}}f(\theta_{B},\theta_{A};x_{i},y_{i}),$ (2)
where $f(\theta_{B},\theta_{A})$ represents the sample-wise loss function. The global loss function is
$F(\theta_{B},\theta_{A})=\sum_{k\in\mathcal{K}^{t}}p_{k}F_{k}(\theta_{B},\theta_{A}),$ (3)
where $p_{k}=|\mathcal{D}_{k}|/|\mathcal{D}|$. To achieve a global model $\theta^{*}=\theta_{P}+\frac{\alpha}{r}\theta_{B}^{*}\theta_{A}^{*}$, clients periodically send updates to the server, assuming no direct interaction among the clients. The optimal global model $\theta^{*}$ for LoRA is identified through the following minimization process
$\theta^{*}=\underset{\theta_{B}\in\mathbb{R}^{d\times r},\theta_{A}\in\mathbb{R}^{r\times\ell}}{\arg\min}F(\theta_{B},\theta_{A}).$ (4)
Within the wireless federated LoRA fine-tuning framework, the objective function $F(\theta_{B},\theta_{A})$ is progressively minimized through a series of local and global updates.
@@
    Broadcasting: The server distributes the $t$-th round global matrices $\theta_{B}^{t}$ and $\theta_{A}^{t}$ to the set of scheduled clients $\mathcal{K}^{t}$ by broadcasting. Since the server has enough power to transmit the global matrices correctly, we assume the downlink transmission is error-free.
@@
    Local Fine-Tuning: Each client $k\in\mathcal{K}^{t}$ initializes the global matrices as an initial point of local matrices, i.e. $\theta_{B,k}^{t}=\theta_{B}^{t},\theta_{A,k}^{t}=\theta_{A}^{t}$. Each client performs local training through gradient descent or ADAM resulting in an updated local models $\theta_{B,k}^{t+1}$ and $\theta_{A,k}^{t+1}$.
@@
where $p_{k}=|\mathcal{D}_{k}|/|\mathcal{D}|$ denotes the proportion of the dataset associated with client $k$, i.e. $\sum_{k\in\mathcal{N}}p_{k}=1$. After a total of $T$ rounds of this process, the training completes, yielding an optimally trained global model for LoRA. Although a client may intend to send the complete matrix product $\theta_{B,k}\theta_{A,k}$, to avoid the computational burden of matrix multiplication at the resource-constrained clients, it instead sends $\theta_{B,k}$ and $\theta_{A,k}$. This is because transmitting the full product would involve sending a matrix of size $d\times\ell$, which is generally much larger than $\theta_{B,k}$ and $\theta_{A,k}$ due to $r\ll\min\{d,\ell\}$. Therefore sending the two low-rank matrices not only minimizes local computation but also reduces the overall communication latency and overhead, as the total number of transmitted parameters is significantly low. Fig. 1 provides an overall framework of the wireless federated LoRA fine-tuning.
@@ II-C Communication Model
We consider a wireless communication model between local clients and the server. In practice, the uplink, which is used by clients to transmit their LoRA matrices to the server, typically experiences more stringent throughput limitations than the downlink. Therefore, our primary focus is on the uplink communication. In contrast, the downlink benefits from higher transmit power, and thus we assume it provides reliable communication at effectively unlimited rates [5]. For the uplink communication, we consider a wireless multiple access system, specifically a frequency division multiple access (FDMA) scheme for local model transmission, with a total available bandwidth of $B$. The channel gain between the server and client is modeled as $h_{k}^{t}=s_{k}^{t}q_{k}^{-\gamma}$, where $s_{k}^{t}\sim\mathcal{N}(0,1)$ represents the small-scale channel fading coefficient, and $q_{k}^{-\gamma}$ denotes the distance-dependent path loss with exponent $\gamma$. Additionally, $b_{k}^{t}\in[0,1]$ represents the bandwidth allocation ratio for client $k\in\mathcal{N}$ where $\sum_{k=1}^{N}b_{k}^{t}=1$. The variable $a_{k}^{t}\in\{0,1\}$ indicates whether client $k$ is selected in iteration $t$, where $\sum_{k=1}^{N}a_{k}^{t}=K^{t}$ and $a_{k}^{t}=0$ implies no bandwidth is allocated to client $k$, resulting in $b_{k}^{t}=0$. The transmission rate for client $k$ can be expressed as $r_{k}^{t}=b_{k}^{t}B\log_{2}\left(1+\frac{|h_{k}^{t}|^{2}}{\sigma^{2}}\right)$, where $\sigma^{2}$ represents the variance of the additive white Gaussian noise (AWGN). During the uplink phase, we assume that the amount of information per model parameter is denoted by $v$. Thus, the communication delay for client $k$ at iteration $t$ can be written as
$D_{k}^{t}=\frac{a_{k}^{t}vr(d+\ell)}{r_{k}^{t}}=\frac{a_{k}^{t}vr(d+\ell)}{b_{k}^{t}B\log_{2}\left(1+\frac{|h_{k}^{t}|^{2}}{\sigma^{2}}\right)}.$ (6)
Note that if $a_{k}^{t}=0$, then $D_{k}^{t}=0$. Since the overall transmission delay at iteration $t$ is determined by the slowest client, it is given by $D^{t}=\max_{k\in\mathcal{K}^{t}}D_{k}^{t}.$